OpenStudy (anonymous):

limit as x -> infinity of (1+1/x)^(4x)

OpenStudy (anonymous):

\[ \lim_{x\to \infty} \left(1+\frac1x\right)^{4x} = \left[\lim_{x\to \infty} \left(1+\frac1x\right)^x\right]^4 = e^4 \]

OpenStudy (anonymous):

@wio how do you know that lim x-> infinity of (1+1/x)^x is e

OpenStudy (anonymous):

It is the definition of \(e\)

OpenStudy (anonymous):

LOLNERD

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

At least one of them.